OpenStudy (anonymous):

I have 2 questions. 13 + 13x and 5/7 - 5/51

OpenStudy (anonymous):

and what is the question :?

OpenStudy (anonymous):

I'm not sure exactly what to do. Trying to help my sister but I'm lost.

OpenStudy (anonymous):

lol,try to read the head question :D

OpenStudy (anonymous):

how old is your sister ?

OpenStudy (anonymous):

middle school

OpenStudy (anonymous):

is it like 13+3x=5/7-5/51 ?

OpenStudy (anonymous):

the first one is this: Use the distributive law to factor each of the following. Check by multiplying. 13 + 13x

OpenStudy (anonymous):

oh ! then you should be aware of this law , then only u'll find the ans. Sorry i dont rem it ! :p

OpenStudy (anonymous):

that's ok. Thank you

OpenStudy (anonymous):

wait beatle ,i will get someone who know it :)

OpenStudy (anonymous):

distributive law r of two types 1. distributive law of multiplication over addition a*(b+c)=(a*b)+(a*c) 2.distributive law of addition over multiplication a+(b*c)=(a+b)*(a+c) it will become 13+13x=(13+13)*(13*x)

OpenStudy (anonymous):

which standard question is this ?

OpenStudy (anonymous):

I don't know what you mean standard? But thank you!

OpenStudy (anonymous):

no problem i was asking that in which class r u?

OpenStudy (anonymous):

oh it's not mine. It's to help my sister.

OpenStudy (anonymous):

ok to solve such questions you must have good concepts of basic properties of mathematics

OpenStudy (anonymous):

I haven't done math in years! haha! I am just terrible but thought to ask to help her.

OpenStudy (anonymous):

thats really good ... by the way what r you studying???

OpenStudy (anonymous):

computers. Just getting started again.

OpenStudy (anonymous):

hay i m too doing graduation in computer sciences

OpenStudy (anonymous):

that's awesome! Thank you for all your help! I appreciate it!

OpenStudy (anonymous):

awww dont mention its wht i do!